OpenStudy (lgbasallote):

an inquisitive physics student and mountain climber climbs a 50 m high cliff that overhangs a calm pool of water. he throws two stones vertically downward, 1 s apart and observes that they cause a single splash. the first stone has an initial speed of 2 m/s. How long after release of the first stone do the two stones hit the water?

OpenStudy (anonymous):

initial velocity=v0=2m/s displacement=h-h0=-50m acc due to gravity=g=9.8m/s^2 time taken by the stones to hit the water=t Calculate 't' using h-h0=(v0)t-(1/2)gt^2.

OpenStudy (lgbasallote):

oh lol of course..what about the initial velocity of the second stone if the two stones hit the water simultaneously? i would equate the distances right?

OpenStudy (lgbasallote):

or do i equate the time?

OpenStudy (anonymous):

both.

OpenStudy (lgbasallote):

i see i see thanks :) i guess i can go from here on

OpenStudy (anonymous):

yw.